Magnetic Corporation expects dividends to grow at a rate of 16.00% for the next two years. After two years dividends are expected to grow at a constant rate of 03.10% indefinitely. Magnetic’s required rate of return is 08.48% and they paid a $2.57 dividend today. Find the value of Magnetic Corporation’s common stock per share by computing:
       a) Dividend at the end of Year 1:
       b) Dividend at the end of Year 2:
       c) Dividend at the end of Year 3:
       d) Price of stock at end of year 2:
       e) Price of stock today:
Answer #1

Current Dividend = D0 = $2.57

Question a:

Dividend at the end of Year 1 = D1 = Do * (1+g) = $2.57 * (1+16%) = $2.9812 = $2.98

Question b:

Dividend at the end of Year 2 = D2 = D1 * (1+g) = $2.9812 * (1+16%) = $3.458192 = $3.46

Question c:

Dividend at the end of Year 3 = D3 = D2 * (1+g) = $3.258192 * (1+3.1%) = $3.56539595 = $3.57

Question d:

Required return =r= 8.48%

Growth rate = g =3.10%

Price of stock at the end of Year 2 = Dividend at the end of year 3 / (Required return - Growth rate)

= $3.56539595 / (8.48% - 3.10%)

= $66.2713002

Price of stock at the end of Year 2 = $66.27

Question e:

Price of Stock today = [D1 / (1+r)^1] + [D2 / (1+r)^2] + [Price of stock at the end of year 2 / (1+r)^2]

= [$2.9812 / (1+8.48)^1] + [$3.358192 / (1+8.48%)^2] +[$66.595092/ (1+8.48%)^2]

= [$2.9812 / 1.0848] + [$3.358192 / 1.17679104] + [$66.595092 /1.17679104]

= $2.74815634 + $2.8536859 + $56.5904139

=62.19

Therefore, price of stock = $62